LCG Experiment Coordination Meeting

→ Europe/Zurich
28-R-015

28-R-015

Description
Coordination meeting between LCG service and LHC experiments, focussing on issues for the coming weeks / months.

Subscribe to lcg-experiment-coordination@cern.ch to be kept informed...

To join the call, do one of the following:
a. Dial +41227676000 (Main) and enter access code 0166222, or
b. To have the system call you, click here: https://audioconf.cern.ch/call/0166222

(Leader code is 0166111)

    • 14:30 → 14:40
      Status of WLCG Q1 Tx-Ty Transfer Milestones - aka Multi-VO tests 10m
      1. Demonstrate Tier0-Tier1 data export at 65% of full nominal rates per site using experiment-driven transfers

        Mixture of disk / tape endpoints as defined by experiment computing models, i.e. ~30% tape for ATLAS; transfers driven by experiments

        Period of at least one week; daily VO-averages may vary (~normal)

      2. Demonstrate Tier0-Tier1 data export at 50% of full nominal rates (as above) in conjunction with T1-T1 / T1-T2 transfers

        Inter-Tier transfer targets taken from ATLAS DDM tests / CSA06 targets

      3. Demonstrate Tier0-Tier1 data export at 35% of full nominal rates (as above) in conjunction with T1-T1 / T1-T2 transfers and Grid production at Tier1s

        Each file transferred is read at least once by a Grid job

        Some explicit targets for WMS at each Tier1 need to be derived from above

    • 14:40 → 14:50
      Status of Dress Rehearsal Preparations - Deployment of Required Services 10m
      • LFC/DPM 1.6.3 is now available. New in LFC: bulk methods for Atlas. New in DPM: production SRM 2.2 support.
      • FTS pilot service is installed with FTS 2.0 and has been tested with low-level dteam transfers - now starting to test its integration with experiment software frameworks. There will be an update on this at tomorrow's Management Board
      • SRM 2.2 tests with FTS continue - starting higher throughput stress-tests to DPM instance at CERN and dCache instance at Fermilab.
      • DB services - see tomorrow's Management Board
    • 14:50 → 15:10
      Feedback from FTS 2.0 pilot service 20m
    • 15:10 → 15:20
      Up-coming service interventions 10m

      2 April is the proposed date for an intervention on 3 network switches and to move the CASTOR nameservers to new hardware. The total downtime is estimated to be 2 hours.

        Goal is to make as many interventions as possible "transparent", i.e. 10' or less and hidden (at least to batch) at middleware and / or experiment frameworks.

        (10' is the time that all systems in the Computer Centre are protected against power glitches).

        Some longer interventions will still be needed, but these too should be minimised:

        • In number;
        • In duration;
        • In impact (i.e. possible service degradation rather than complete outtage).

        Possible slots prior to start of Dress Rehearsals and Pilot LHC run:

        • Weeks before and / or after Easter (weeks of 2 and 9 April)
        • Slot in June, e.g. June 11 (following weeks are CMS / ATLAS weeks respectively).
        • Week of CHEP 2007?? (1st week of September)

        Foreseen interventions:

        • Change of network switches affecting numerous Grid services

          MID-RANGE: Computer center (513-C): IP301 to 303

          To check the machines behind these services:

          https://network.cern.ch/sc/fcgi/sc.fcgi?Action=ListStarpoints

          CS estimate around 2 hours per service, depending on the number of switches (1 to 4)which means that all machines behind that switch are off the network for 2 hours.

        • Move of CASTOR DBs to better h/w (to be done by VO - service interruption involved).
        • Deployment of new CASTOR LSF plug-in
        • CASTOR disk server h/w fixes (or rolling replacement of servers...)
        • Oracle Critical Patch updates
          • 17 April 2007
          • 17 July 2007
          • ???16 October 2007???
          • 15 January 2008
    • 15:20 → 15:40
      Main activities for Q2 2007 20m
    • 15:40 → 15:45
      AOB 5m